If a right triangle has two legs of 15 ft and 18 ft what is the hypotenuse and how do you get the answer?
Answer by nerdybill(7384)

If a right triangle has two legs of 15 ft and 18 ft what is the hypotenuse and how do you get the answer?
Because it is a "right triangle", Pythagorean theorem applies:
a^2 + b^2 = c^2
where
c is the hypotenuse
a and b are the other two sides
.
plug in the given info:
15^2 + 18^2 = c^2
225 + 324 = c^2
549 = c^2
23.43 ft = c (hypotenuse)